# Spatially-Embedded Lens Visualization: A Design Space

### v1

@@ 2.1 Position, Orientation & Scale
Most lens techniques supply users full interactivity for setting the lens position, orientation, and scale.
@@
Although most surveyed lenses are fully interactive, some rely on (semi-)automated control based on the underlying data [18, 1, 3].
@@ 2.2 Shape
Most lenses proposed in the literature are of fixed shape.
@@
Some lens techniques adapt their shape dynamically, either by the user or by the underlying data
@@ 2.3 Dimensionality
2D lenses are the most commonly used and have polygonal shapes—-e.g. circular or square, which are placed and manipulated in screen space.
@@
2.5D lenses consist of 2D lenses embedded in 3D spaces, thereby augmenting lens manipulations to 3D [42, 33, 23, 6]. Although this alleviates the issue of spatial correlation, it increases interaction effort to place and align the lenses according to the underlying data. In addition, the flat nature of 2.5D lenses constraint the spatial selection to only a slice of the 3D data. To illustrate, Spindler et al. introduced a spatially aware, handheld paper sheet lens to improve lens manipulations in 3D spaces above tabletop displays [53]—e.g. users could arbitrarily slice an MRT scan of a human torso in different poses.
3D lenses consist of 3D volumes—e.g. a sphere,
@@ 2.4 Effect Scope
A few lens effects impact visual elements located not only inside the lens boundary but also beyond it; therefore, the effect scope is considered to be the lens exterior [41, 50, 19]. In such cases, the lens boundary is primarily used to define a selection based on a local behavior of interest in the underlying data—e.g. to select trajectories based on a local directionality.
@@
This effect scope is referred to as separate view [46, 10, 50, 27, 58].
@@ 2.5 Effect Imagery
In 2D lens images the imagery generated as the lens effect consists of projected images, which are rendered either directly onto the view plane [23, 24] or onto a separate render texture to display the lens image as a separate view [42, 46, 6, 5].
@@
3D images consist of volumetric visual images, possibly rendered as a separate volume and enriched with additional visual elements [9, 28, 5]—similar to a world-in-miniature [54].
@@
Decal images conform to surface areas of interest. This conforming behavior not only establishes a strong spatial correlation between the lens image and its referent location but also mitigates perceptual issues, such as the lens image itself occluding parts of the base visualization—particularly those of intricate geometry.
@@ 2.6 Effect Encoding
Thematic encoding draws from information visualization and refers to abstract representations of the lens’ data selection—e.g. linear and radial histograms (Fig. 8–left) [58, 40, 36].
@@ 2.7 Viewpoint Dependency
Although most papers do not consider viewpoint in their designs, view-sensitive lenses afford fluid and intuitive interactions by dynamically adapting to users’ physical navigation or gestures—e.g. automatically expanding when users move away, and revealing fine-grained information when in close proximity (Fig. 9–right) [32]. These dynamic responses allow lenses to better anticipate and align with user intent, likely reducing the need for manual input. However, design considerations are needed to avoid unstable and disorienting effects caused by unintentional movements—e.g. to use smoothing or “lock-in” mechanisms.

### v2

@@ 2 Design Space
), Effect Scope (
), Shape (
), Dimensionality (
), Effect Imagery (
), Effect Encoding (
), and Viewpoint Dependency (
) dimensions. Some references contain more than one colored cell within a single dimension, given its non-mutually exclusive design choices. Papers presenting lens techniques within immersive spaces are marked with an asterisk (
@@ 2.1 Position, Orientation & Scale
Most lens techniques supply users full interactivity
for setting the lens position, orientation, and scale.
@@
Although most surveyed lenses are fully interactive, some rely on (semi-)automated
control based on the underlying data [18, 1, 3].
@@ 2.2 Shape
Most lenses proposed in the literature are of fixed
shape.
@@
Some lens techniques adapt their shape dynamically, either by the user
or by the underlying data
@@ 2.3 Dimensionality
2D lenses
are the most commonly used and have polygonal shapes—-e.g. circular or square, which are placed and manipulated in screen space.
@@
2.5D lenses
consist of 2D lenses embedded in 3D spaces, thereby augmenting lens manipulations to 3D [42, 33, 23, 6]. Although this alleviates the issue of spatial correlation, it increases interaction effort to place and align the lenses according to the underlying data. In addition, the flat nature of 2.5D lenses constraint the spatial selection to only a slice of the 3D data. To illustrate, Spindler et al. introduced a spatially aware, handheld paper sheet lens to improve lens manipulations in 3D spaces above tabletop displays [53]—e.g. users could arbitrarily slice an MRT scan of a human torso in different poses.
3D lenses
consist of 3D volumes—e.g. a sphere,
@@ 2.4 Effect Scope
A few lens effects impact visual elements located not only inside the lens boundary but also beyond it; therefore, the effect scope is considered to be the lens exterior
[41, 50, 19]. In such cases, the lens boundary is primarily used to define a selection based on a local behavior of interest in the underlying data—e.g. to select trajectories based on a local directionality.
@@
This effect scope is referred to as separate view
[46, 10, 50, 27, 58].
@@ 2.5 Effect Imagery
In 2D lens images
the imagery generated as the lens effect consists of projected images, which are rendered either directly onto the view plane [23, 24] or onto a separate render texture to display the lens image as a separate view [42, 46, 6, 5].
@@
3D images
consist of volumetric visual images, possibly rendered as a separate volume and enriched with additional visual elements [9, 28, 5]—similar to a world-in-miniature [54].
@@
Decal images
conform to surface areas of interest. This conforming behavior not only establishes a strong spatial correlation between the lens image and its referent location but also mitigates perceptual issues, such as the lens image itself occluding parts of the base visualization—particularly those of intricate geometry.
@@ 2.6 Effect Encoding
Thematic encoding
draws from information visualization and refers to abstract representations of the lens’ data selection—e.g. linear and radial histograms (Fig. 8–left) [58, 40, 36].
@@ 2.7 Viewpoint Dependency
Although most papers do not consider viewpoint in their designs, view-sensitive
lenses afford fluid and intuitive interactions by dynamically adapting to users’ physical navigation or gestures—e.g. automatically expanding when users move away, and revealing fine-grained information when in close proximity (Fig. 9–right) [32]. These dynamic responses allow lenses to better anticipate and align with user intent, likely reducing the need for manual input. However, design considerations are needed to avoid unstable and disorienting effects caused by unintentional movements—e.g. to use smoothing or “lock-in” mechanisms.